OpenStudy (nuccioreggie):

Dan and Barney are members of different video game libraries. Dan pays a membership fee of $35, and he pays $4 for every video game he rents. The following function shows the total amount of money, y, in dollars, that Barney pays for renting x number of video games: y = 3x + 30. How many more dollars does Dan pay for a membership fee than Barney? $1 $5 $26 $32

OpenStudy (mathmath333):

the membership fee of barney can be calculated when x=0 in below equation, y = 3x + 30. find value of y first when x=0

OpenStudy (nuccioreggie):

3

OpenStudy (mathmath333):

y=3*0+30= ?

OpenStudy (nuccioreggie):

30

OpenStudy (mathmath333):

the more dollars that Dan pays = 35-30= ?

OpenStudy (nuccioreggie):

5